QUESTION 34
We are analyzing the effects of regime type on corruption rates with the following model: Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy where Corruption is an index of corruption, GDP (per capita) is measured in thousands of dollars, and Democracy is a dummy variable that is equal to one if a country is a democracy and 0 otherwise. What is the expected rate of corruption of a democratic country with a per capita GDP of $50,000?
   A.   8.5
   B.   Not enough information to come up with an answer
   C.   13.0
   D.   - 4992
2.5 points   

QUESTION 35
We are analyzing the effects of regime type on corruption rates with the following model: Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy where Corruption is an index of corruption, GDP (per capita) is measured in thousands of dollars, and Democracy is a dummy variable that is equal to one if a country is a democracy and 0 otherwise. Suppose we want to know the estimated effect on corruption of an extra thousand dollars per capita for a democratic country. Our estimate implies the change in predicted corruption will be
   A.   0.1 higher
   B.   2.1 higher
   C.   2.1 lower
   D.   0.1 lower
2.5 points   

QUESTION 36
We are analyzing the effects of regime type on corruption rates with the following model: Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy where Corruption is an index of corruption, GDP (per capita) is measured in thousands of dollars, and Democracy is a dummy variable that is equal to one if a country is a democracy and 0 otherwise. What is the expected rate of corruption of a democratic country with a per capita GDP of $50,000?
   A.   13.0
   B.   Not enough information to come up with an answer
   C.   8.5
   D.   - 4992
2.5 points   

QUESTION 37
When dealing with categorical variables in the context of a multivariate regression, we:
   A.   Change the categorical variable into a dummy variable (per category) and include all of the dummy variables in the regression.
   B.   Include the categorical variable directly into the model and interpret the results
   C.   Interact the categorical variable with a continuous variable.
   D.   Change the categorical variable into a dummy variable (per category) and include all but one of the dummy variables in the regression.
2.5 points   

QUESTION 38
Which of the following are consequences of measurement error in the dependent variable?
   A.   Smaller R2
   B.   Biased coefficient estimates
   C.   The bigger the measurement error, the bigger the variance of the error term.
   D.   Both A and C
2.5 points   

QUESTION 39
Which of the following is NOT a categorical variable:
   A.   Car manufacturer
   B.   Level of Education
   C.   Region
   D.   Race
2.5 points   

QUESTION 40
Which of the following will tend to reduce the size of a confidence interval?
   A.   Decrease the standard deviation of the population
   B.   Decrease the significance level (alpha).
   C.   You can't do anything to reduce the interval
   D.   Increase the sample size
Answer #1

Statistics and Probability

QUESTION 34
We are analyzing the effects of regime type on corruption rates with the following model: Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy where Corruption is an index of corruption, GDP (per capita) is measured in thousands of dollars, and Democracy is a dummy variable that is equal to one if a country is a democracy and 0 otherwise. What is the expected rate of corruption of a democratic country with a per capita GDP of $50,000?
   A.   8.5
   B.   Not enough information to come up with an answer
   C.   13.0
   D.   – 4992

2.5 points   

Explanation:

The fitted model is

Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy

Using above model, the expected rate of corruption of a democratic country with a per capita GDP of $50,000 is calculated as

Corruption = 10 - 0.1*50,000- 2.0(1)=-4992

QUESTION 35
We are analyzing the effects of regime type on corruption rates with the following model: Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy where Corruption is an index of corruption, GDP (per capita) is measured in thousands of dollars, and Democracy is a dummy variable that is equal to one if a country is a democracy and 0 otherwise. Suppose we want to know the estimated effect on corruption of an extra thousand dollars per capita for a democratic country. Our estimate implies the change in predicted corruption will be


   A.   0.1 higher
   B.   2.1 higher
   C.   2.1 lower
   D.   0.1 lower

2.5 points   

Explanation: We have coefficient of GDP (per capita) =-0.1, thus unit change in GDP (per capita) will bring Corruption down by 0.1 times.

QUESTION 36
We are analyzing the effects of regime type on corruption rates with the following model: Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy where Corruption is an index of corruption, GDP (per capita) is measured in thousands of dollars, and Democracy is a dummy variable that is equal to one if a country is a democracy and 0 otherwise. What is the expected rate of corruption of a democratic country with a per capita GDP of $50,000?
   A.   13.0
   B.   Not enough information to come up with an answer
   C.   8.5
   D.   – 4992


2.5 points

Explanation:

The fitted model is

Corruption = 10 - 0.1 GDP (per capita) - 2.0Democracy

Using above model, the expected rate of corruption of a democratic country with a per capita GDP of $50,000 is calculated as

Corruption = 10 - 0.1*50,000- 2.0(1)=-4992  

QUESTION 37
When dealing with categorical variables in the context of a multivariate regression, we:
   A.   Change the categorical variable into a dummy variable (per category) and include all of the dummy variables in the regression.
   B.   Include the categorical variable directly into the model and interpret the results
   C.   Interact the categorical variable with a continuous variable.
   D.   Change the categorical variable into a dummy variable (per category) and include all but one of the dummy variables in the regression.
2.5 points

Answer: D.   Change the categorical variable into a dummy variable (per category) and include all but one of the dummy variables in the regression.

QUESTION 38
Which of the following are consequences of measurement error in the dependent variable?
   A.   Smaller R2
   B.   Biased coefficient estimates
   C.   The bigger the measurement error, the bigger the variance of the error term.
   D.   Both A and C
2.5 points

Answer: D.   Both A and C

QUESTION 39
Which of the following is NOT a categorical variable:
   A.   Car manufacturer
   B.   Level of Education
   C.   Region
   D.   Race
2.5 points   

Answer: A.   Car manufacturer

QUESTION 40
Which of the following will tend to reduce the size of a confidence interval?
   A.   Decrease the standard deviation of the population
   B.   Decrease the significance level (alpha).
   C.   You can't do anything to reduce the interval
   D.   Increase the sample size

Answer: A.   Decrease the standard deviation of the population